# CAPRI-FAIR: Integration of Multi-sided Fairness in Contextual POI Recommendation Framework

### v1

@@ 5. Conclusion
In order to integrate consumer and provider fairness into pre-existing POI recommendation models, a post-filter approach was employed using 2 fairness factors added into the pre-filter score with varying weights. The provider factor gives higher scores for less popular POIs, using a linear model fit to the popularity distribution, such that recommendations with moderate popularity are not severely penalized. The consumer factor attempted to increase recommendation quality for inactive users by recommending more popular POIs near previously visited ones.
The results show that increasing the provider factor’s weight resulted in improved exposure to the long-tail, up to 3-fold in the Yelp dataset, especially for USG where exposure was almost none. This came at the cost of lower precision, varying between models. The consumer factor improved precision in some cases but failed to enhance performance for inactive users. Nevertheless, combining both factors can yield reasonable results. Overall, there is a tradeoff between these factors, with the provider fairness factor achieving better long-tail exposure with minimal precision decreases for models like USG. Unfortunately, when comparing GCE metrics across models, a distinct correlation exists between low item exposure GCE values and higher precision, although this depends on the dataset, as the LORE model reached a middle ground between the metrics in the Gowalla dataset.
In the future, other models and strategies for consumer fairness can be assessed, and other objectives such as system stakeholder metrics can be incorporated in an attempt to balance multiple objectives.
@@ A.2. Modelling the provider fairness factor using various functions
A.3. More results on the tradeoff between provider and consumer fairness

### v2

@@ 5. Conclusion
To integrate consumer and provider fairness into pre-existing POI recommendation models, a post-filter approach was employed using 2 fairness factors added into the pre-filter score with varying weights. The provider factor gives higher scores for less popular POIs, using a linear model fit to the popularity distribution, so moderately popular recommendations are not severely penalized. The consumer factor attempted to increase recommendation quality for inactive users by recommending more popular POIs near previously visited ones. The results show that increasing the provider factor’s weight resulted in improved exposure to the long-tail, up to 3-fold in the Yelp dataset, especially for USG where exposure was almost none. This came at the cost of lower precision, varying between models. In some cases, the consumer factor improved precision but failed to enhance performance for inactive users. Nevertheless, combining both factors can yield reasonable results. Overall, there is a tradeoff between these factors, with the provider fairness factor achieving better long-tail exposure with minimal precision decreases for models like USG. Unfortunately, when comparing GCE metrics across models, a distinct correlation exists between low item exposure GCE values and higher precision, although this depends on the dataset, as the LORE model reached a middle ground between the metrics in the Gowalla dataset.
In the future, we plan to expand our research to include various models and strategies for consumer fairness and other objectives. We will explore different splitting methods and refine our accuracy assessments, aiming to comprehensively address complex fairness issues and enhance POI recommender systems’ effectiveness.
ACKNOWLEDGMENTS
This research is partly funded by the Australian Research Council Centre of Excellence for Automated Decision-Making and Society (ADM+S) (CE200100005).
@@ A.2. Modelling the provider fairness factor using various functions
A.3. Mann-Whitney U tests for the provider fairness scoring models
Aside from the Kruskal-Wallis tests ran across the 3 scoring models, we also ran pairwise Mann-Whitney U tests. The significance results are shown below in Figure 6, referencing the values in Table 1. Note that the differences in long-term exposure are immediately significant even for low factor weights. The significance in differences in precision come a little bit later with higher factor weights, specifically in the Yelp dataset.
Yelp Dataset Precision L-T Exposure
Model Comparison 0.25 0.50 0.75 1 0.25 0.50 0.75 1
GeoSoCa Lin. v.s. Power.
Power. v.s. Log.
Lin. v.s. Log.
LORE Lin. v.s. Power.
Power. v.s. Log.
Lin. v.s. Log.
USG Lin. v.s. Power.
Power. v.s. Log.
Lin. v.s. Log.
Gowalla Dataset Precision L-T Exposure
Model Comparison 0.25 0.50 0.75 1 0.25 0.50 0.75 1
GeoSoCa Lin. v.s. Power.
Power. v.s. Log.
Lin. v.s. Log.
LORE Lin. v.s. Power.
Power. v.s. Log.
Lin. v.s. Log.
USG Lin. v.s. Power.
Power. v.s. Log.
Lin. v.s. Log.
Figure 6. Significance results of pairwise Mann-Whitney U tests for different provider fairness scoring models.
A.4. More results on the tradeoff between provider and consumer fairness